Circular No. 6124

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET 1995 A1 Robert Jedicke, Lunar and Planetary Laboratory, reports his discovery

of a comet with the 0.9-m Spacewatch telescope at Kitt Peak. J. V. Scotti adds that the middle image gives m_{2} = 20.9, shows a coma of diameter 17′′ and a tail extending faintly out to 2′.91 in p.a. 262^{o}.

1995 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 8.41727  7^{h}56^{m}12.^{s}11  +20^{o}20′54′′.9  19.0
8.46370  7 56 10.50  +20 21 06.1
8.49853  7 56 09.37  +20 21 15.0

NOTICE REGARDING NAMES AND DESIGNATIONS OF COMETS

In accordance with the IAU resolution on MPC 23803–23804 (summarized on IAUC 6076), the name of comet 1995 A1 will be established in consultation with the Commission 20 Small Bodies Names Committee and will not be announced until a preliminary orbit is available. For convenience, we give here the correspondences between old-style and new-style designations of comets discovered during the past six months: 1994m = C/1994 N1 (Nakamura-Nishimura-Machholz), 1994n=P/1994 N2 (McNaught-Hartley), 1994o = P/1994 P1 (Machholz 2), 1994r = C/1994 T1 (Machholz) and 1994u = P/1994 X1 (McNaught-Russell). In general, there will not be new-style designations for comet recoveries, and use will be made instead of the sequential numbers given for such comets on MPC 24252 (e.g., 73P/Schwassmann-Wachmann 3 for the recovery 1994w). The first-return comet recovery 1994v = 1990a = 1990 X becomes P/1994 V1 (Wild 4) and will receive a sequential number shortly.

SUPERNOVA 1994ak IN NGC 2782 P. Garnavich, P. Challis, A. Riess and R. Kirshner, Center for Astro-

physics, report that a spectrum of SN 1994ak was obtained by P. Berlind on 1994 Dec. 28.45 UT with the 1.5-m Tillinghast Telescope on Mt. Hopkins. The spectrum shows strong, narrow (FWHM 400 km/s) Balmer emission lines near the redshift of NGC 2782, superposed on a featureless blue continuum. He I (587.5 nm and 667.8 nm) is also present. The base of the lines is broad (full width at the continuum of 5000 km/s), and a weak P-Cyg absorption appears with the minimum at –600 km/s. The estimated absolute magnitude of this object implies that it is a supernova. The Balmer emission is similar to that observed in SN 1988Z, and 1994ak is therefore likely to be “type IIn” (Schlegel 1990, MNRAS 244, 269) at an early phase.

1995 January 9 Brian G. Marsden